# Variational formulations of transport on combinatorial meshes

### v6

@@ Abstract
We develop analogues of the primal and mixed variational formulations of the conservation laws describing transport of scalar physical properties for topological spaces referred to as cell complexes. The development is restricted to cell complexes with cells having connectivity of simple polytopes. Such spaces are suitable representations of physical systems composed of elements of different apparent topological dimensions, where all elements might have individual properties, and elements of a given dimension might interact via common elements of lower dimensions. This modelling foundation can be considered as intermediate between the foundation for particle-based modelling, which is the discrete topology, and the foundation for continuum modelling, which is the smooth topology. The new foundation offers advantages for the analysis of the behaviour of physical systems with complex internal structures. The development is based on a calculus with combinatorial differential forms which is a discrete analogue of the smooth exterior calculus with differential forms. We call the resulting calculus combinatorial mesh calculus (CMC) as it is based on combinatorial meshes, for which embedding is forgotten and only the connectivity between cells and measures of the cells are used in calculations. We discuss how the obtained formulation is specialised for several different problems, including mass diffusion, heat conduction, fluid flow in porous media, and charge diffusion, and provide details about the formulation of initial boundary value problems for these transport cases. Examples and results for selected boundary value problems are given to demonstrate the capabilities of the CMC formulations.
@@ 1.1 Motivation
Modern experimental techniques are revealing, with increasing details, the complexity of the internal structures of materials [1, 2] and the evolution of these structures caused by various forces [3, 4]. The observations clearly show that the internal structures on any length scale consist of finite-size components that appear with different dimensions. For example, relative to the bulk on a given length scale, monolithic materials contain defects, and composites contain intentionally inserted inclusions that appear as points, lines, and surfaces. Specifically, at the polycrystalline level, metals and alloys appear to be collections of 3D (grains), 2D (grain or interphase boundaries) and 1D (boundary junctions). At sub-grain level, they appear to have 3D (regions with perfect atomic lattice), 2D (stacking faults), and 1D (dislocations) components. This view can be extended to the atomic scale, where the 3D components are the Wigner-Seitz cells around atoms, while the 2D and 1D components are their faces and edges [5].
The behaviour of a structure at a given length scale emerges from interactions between components of all dimensions. Assuming local interactions, this means that in principle any given component affects all components of lower dimensions on its boundary, as well as all components of higher dimensions which have it on their boundaries. The observations also clearly show that the evolution of a structure at a given length scale is a sequence of discrete, finite changes of the arrangement, and/or shape, and/or size, and/or nature of its components due to, e.g., shear, separation, diffusion, dissolution, chemical/electrochemical reactions, etc. The components of different dimensions and nature react to and transmit differently the evolution forces, i.e., the internal structures localise and channel the evolution.
The understandings of the finiteness of microstructural components and evolution events, as well as the geometry of local interactions, are not consistently reflected in the existing methods for materials’ modelling. The state-of-the-art offers advanced representations of observed structures [6, 7, 8, 9] but the descriptions of interactions and evolution mechanisms do not respect all observations because of the constraints imposed by their topological foundations. The descriptions used to date have been developed for two conceivable topologies: a discrete topology, $T(D)$, where the system is considered as a collection of finitely many distinct constituents (points or particles); and a continuous/smooth topology, $T(C)$, where the system is considered as a collection of infinitely many infinitesimal points forming a smooth manifold.
$T(D)$ is the foundation for several modelling approaches, most popular of which are molecular dynamics [10], peridynamics [11], smoothed particle hydrodynamics [12], and discrete element method [13]. $T(D)$ dictates operations with discrete fields over constituents or pairs of constituents, namely discretely distributed extensive physical properties (mass, energy, momenta, charge), discretely distributed fluxes of these properties, and discretely distributed material characteristics (diffusivity, conductivity, stiffness, etc.). The resulting methods are useful for capturing atomic and mesoscopic phenomena but their representation of material structures lacks explicit topological connectivity, making it difficult to enforce constraints and predict emergent material behaviours accurately. Furthermore, their lack of an intrinsic notion of volume means that they struggle to handle long-range interactions effectively. A significant drawback of $T(D)$-based methods is the necessity for heuristic force laws, which may not always be physically justified.
@@
$T(C)$ is the foundation for continuum physics and mechanics [14]. These operate with fields of intensive (averaged) physical properties, which are densities of the extensive properties. This foundation leads to formulations of the conservation laws in integral (weak) form and, under the assumption of locally continuous fields, in differential (strong) form. With very few exceptions with specific domain geometries and boundary conditions, the continuum formulations have no closed-form solutions and require approximate numerical approaches. The most popular are the finite element methods [15], the finite difference methods [16], and the finite volume methods [17]. These methods effectively capture macroscopic mechanical properties but struggle with representing discontinuities such as cracks, defects, and phase boundaries. These are presently considered by different numerical approximations, such as cohesive zone models [18, 19], extended finite element method [20, 21], and phase-field methods [22, 23]. The latter use artificial smoothing to handle sharp transitions, which can lead to inaccuracies in representing localized phenomena. Finally, continuum-based methods rely on constitutive laws that may not always adequately capture the complexities of real-world materials, particularly in cases involving extreme deformations, multi-phase interactions, or non-equilibrium processes.
@@
The constraints of the discrete and continuous topologies can be avoided by adopting another topological foundation for modelling - cell complexes. These are collections of cells of different topological dimensions arranged/glued following specific rules and are studied in algebraic topology [24, 25]. Unlike the continuum approach, which relies on smooth differentiable manifolds, or discrete particle methods that lack explicit topological structure, cell complexes naturally encode the connectivity and geometry of complex material domains. This makes them particularly well-suited for modelling materials with intricate microstructures, such as polycrystalline metals, composite materials, and biological tissues. Therefore, we will refer to cell complexes as the natural topology $T(N)$.
@@
In a previous work [26], we showed a formulation of the conservation of scalar quantities (mass, energy, charge) in cell complexes that was analogous to the strong (differential) form in the continuum settings. The formulation used combinatorial differential forms, introduced by Forman [27], for which we proposed the metric-dependent operations required to describe the physics, showed the construction of boundary value problems on cell complexes, and applied the method to analysis of diffusion in composites with complex arrangements of components. One drawback of the strong form’s analogue is the approximate way of applying Neumann boundary conditions. The weak (integral) formulations of the conservation laws avoid this approximation.
1.2 Contribution
The aim of this work is to develop the two weak formulations of conservation of scalar quantities on cell complexes – primal (in terms of the scalar potential only), and mixed (in terms of the scalar potential and the quantity flow rate). The weak formulations with combinatorial differential forms are analogous to the continuum weak formulations with exterior differential forms. The latter have not been published to date to the best of our knowledge. It is therefore necessary and instructive to make this contribution first before describing the discrete analogues. This also helps to identify the similarities and differences between the continuum and discrete operations and formulations. The continuum formulations are derived directly with exterior differential forms, without reference to the traditional vector calculus. To this end, we provide an overview of smooth exterior calculus in Section 2 and then derive the continuum formulations in Section 3. The basic notions and operations related to cell complexes required for this work will be introduced in Section 4. These are then used to describe the physics on cell complexes in Section 5, with rigorous description of the boundary conditions for (initial) boundary value problems. The $T(N)$-based formulations are used to solve a series of boundary value problems with domain geometries and boundary conditions that have closed form (exact) continuum solutions to facilitate comparison. Selected examples are given in Section 6, and a larger set of examples can be found at https://github.com/kipiberbatov/cmc.
@@ 1.2 Contribution
The sections are broken down into items named discussion, definition, notation, remark, proposition, corollary, formulation, and example, which are numbered consecutively to facilitate cross-referencing. All unary operators (excluding taking the additive inverse) have higher precedence than binary operators (excluding taking exponents). For instance, $fv\wedge w$ will mean $(fv)\wedge w$. Further, we record extensively the physical dimensions of the quantities, operators, and equations we discuss. If a quantity $v$ has dimension $X$, we will write $v[X]$ when we want to record the dimension of $v$, and $[[v]]=X$ when we do calculations with dimensions. Physical dimensions will be represented by the SI quantities for mass $\mathsf{M}$, length $\mathsf{L}$, time $\mathsf{T}$, temperature $\theta$, and charge $\mathsf{C}$. (All except the last one are base SI quantities; for historical reasons electric current $\mathsf{I}=\mathsf{C}\mathsf{T}^{-1}$ is used instead of $\mathsf{C}$ as a base quantity, but we prefer charge since it is a conserved property of matter.) If an operator $\varphi$ takes quantities $v_{1}[X_{1}],...,v_{n}[X_{n}]$ and returns a quantity $\varphi(v_{1},...,v_{n})[YX_{1}...X_{n}]$, we will say that $\varphi$ is of dimension $Y$, and write $\varphi[Y]$ or $[[\varphi]]=Y$. Finally, for equations, if $v[X]$ and $w[X]$, we will emphasise their dimensions in equalities, writing $v=w[X]$.
@@ 2 Overview of exterior calculus
In this section, we provide an overview of the exterior calculus apparatus required in the rest of the work. Some basic knowledge of general topology is required. Experience with smooth manifolds is also beneficial, although we introduce the terminology we use anyway. A comprehensive book on smooth manifolds is, for instance, [28]. We find many benefits in using exterior calculus instead of vector calculus.
@@
    There is a clear distinction between manifold’s intrinsic (non-metric) calculus (Section 2.1) and (additional) calculus depending on a structure such as a Riemannian metric (Section 2.2). This elucidates the distinction between different types of physical relations and the mapping of these relations straightforwardly to their discrete counterparts, for which we build a similar calculus.
@@
    Physical quantities are represented by differential forms ready to be integrated over their respective domains. In contrast, in $3$D vector calculus, there are only scalar and vector fields, and is impossible to differentiate between “scalar” quantities living on 00- and $3$-cells, and “vector” quantities living on $1$-cells and $2$-cells. Terms like pseudo-scalar fields and pseudo-vector fields are often used, but they are related to sign invariance under change of orientation. Because sign invariance is an important notion, we will use of the term pseudo-forms.
@@ 2.1 Topological operations
It can be proven [28, Corollary 3.3] that for any $y\in M$, $T_{y}M$ is a $D$-dimensional real vector space, and $\mathfrak{X}M$ is a $D$-dimensional module over $\mathcal{F}M$. In local coordinates (Cartesian coordinates when $M=\mathbb{R}^{D}$) a vector field $X$ and a tangent vector $\left.X\right|_{y}$ have the representation
@@
The dual space $V^{*}$ of $V$ is defined as $\mathop{\rm Hom}\nolimits(V,R)$, the space of linear functions on $V$. If $V$ is finite dimensional, $\dim V=D$, then also $\dim V^{*}=D$. To a basis $(e_{1},...,e_{D})$ of $V$ corresponds the dual basis $(e^{1},...,e^{D})$ of $V$, so that for $1\leq p,q\leq D$, $e^{p}(e_{q})=\delta^{p}_{q}$, where $\delta^{p}_{q}$ is the Kronecker delta ($1$ if $p=q$ and zero if $p\neq q$).
The exterior algebra $\Lambda^{\bullet}V$ of $V$ is the freest associative and alternating algebra containing $V$, i.e., it is the algebra generated by an associative operation $\wedge$, the wedge (exterior) product, subject to the alternating property: for all $v\in V$, $v\wedge v=0$. From the alternating property it follows the antisymmetric property: for any $v,w\in V$, $v\wedge w=-w\wedge v$. When $V$ is of dimension $D$ with basis $e_{1},...,e_{D}$, $\Lambda^{\bullet}V$ is decomposed into
@@
[…]
@@
where the right hand side is the standard derivative of the real-valued function $f\circ\gamma\colon[0,\delta)\to\mathbb{R}$. We say that $v$ is outward-pointing if $-v$ is inward pointing.
@@
is outward-pointing. Indeed, the curve $\gamma(t)=(1-t)(\cos\varphi_{0},\sin\varphi_{0})$ witnesses the fact that $-v$ is inward-pointing.
@@
    $i$ is a pairing on covectors: for any $v\in V$ and $\omega\in\Lambda^{1}V^{*}\simeq V^{*}$,
@@
Let $M$ be an oriented smooth manifold of dimension $D$ with boundary and orientation $\mathop{\rm or}\nolimits_{M}$, and $S$ be an oriented $(D-1)$-dimensional submanifold of $\partial M$ with orientation $\mathop{\rm or}\nolimits_{S}$. Consider an outward-pointing vector field $X$ on $S$. Since $\dim_{\mathcal{F}M}(\Omega^{D-1}S)=1$, there exists unique smooth function $f\colon S\to\mathbb{R}\setminus\{0\}$ such that
@@
Consider an oriented manifold $M$ with boundary $N$, $D=\dim M$, and let $P$ be an oriented $(D-2)$-submanifold of $N$. Locally $P$ divides $N$ into exactly two parts $N^{\prime}$ and $N^{\prime\prime}$. (This is the diamond property – there are exactly two “cells” between $M$ and $P$. A similar property holds for convex polytopes as discussed in [29, Theorem 2.7 (iii)].) Moreover, if $N^{\prime}$ and $N^{\prime\prime}$ have the same orientations as $N$ (the induced boundary orientation), then they will have opposite relative orientations with $P$, i.e., $\varepsilon(N^{\prime},P)=-\varepsilon(N^{\prime\prime},P)$. To allow for possible non-induced orientations on $N^{\prime}$ and $N^{\prime\prime}$, we can write the following equality [26, Supplementary matrial, Proposition Appendix B.13] which is satisfied in general:
@@
[…]
@@
    (base case) integration is function evaluation for 00-forms: for any singleton 00-dimensional manifold $\{x\}$ with orientation $\varepsilon\in\{-1,1\}$, and any $f\colon\{x\}\to\mathbb{R}$,
@@
Let $M$ be an oriented smooth manifold with boundary, $D=\dim M$. An important consequence of the graded Leibniz rule, eq. 2.18, and the Stokes-Cartan theorem, eq. 2.47, is the integration by parts formula: for any $p\in\{0,...,D-1\},\ \omega\in\Omega^{p}_{c}M,\ \eta\in\Omega^{D-p-1}_{c}M$,

### v7

# Variational formulations of transport phenomena on combinatorial meshes

@@ Abstract
We develop primal and mixed variational formulations of transport phenomena on cell complexes with simple polytope connectivity. This framework addresses materials with internal structures comprising components of different topological dimensions, where cells of each dimension may possess distinct physical properties. The approach, which we call Combinatorial Mesh Calculus (CMC), extends Forman’s combinatorial differential forms, previously used to formulate strong conservation laws. CMC operates directly on meshes without requiring smooth embeddings, using discrete analogues of the exterior derivative, Hodge star, and co-differential operators. Our mixed formulation achieves computational efficiency through diagonal stiffness matrix A, which admits direct inversion and enable efficient solution strategies. CMC differs from Discrete Exterior Calculus, which requires circumcentric duality and well-centred meshes, and from Finite Element Exterior Calculus, which constructs polynomial spaces on smooth domains. Our framework applies to general cell complexes, including curved cells and irregular meshes without geometric quality constraints. The mathematical development proceeds in parallel between the smooth and discrete settings, establishing correspondences between continuous and discrete operators. Initial boundary value problems are formulated for mass diffusion, heat conduction, charge transport, and fluid flow through porous media. Numerical examples on regular and irregular meshes in two and three dimensions demonstrate agreement with analytical solutions. The framework enables modelling of transport in materials where microstructural topology influences macroscopic behaviour, with applications to polycrystalline materials, composites, and porous media.
@@ 1.1 Motivation
Modern experimental techniques are revealing, with increasing detail, the complexity of internal structures in materials kumar2023MC , hayashi2024MC and their evolution under various forces stock2020XCT , woodruff2021SCT .
@@
These observations demonstrate that material structures at any length scale consist of finite-size components appearing with different topological dimensions. Relative to the bulk, materials contain point defects, line defects, and surface defects. For example, at the polycrystalline level, metals and alloys appear as collections of 3D grains, 2D grain or interphase boundaries, and 1D boundary junctions. At sub-grain level, they contain regions with perfect atomic lattice (3D), stacking faults (2D), and dislocations (1D). This multi-dimensional view extends to the atomic scale, where 3D Wigner-Seitz cells around atoms connect through 2D faces and 1D edges hunklinger2022SSP .
The behaviour of these structures emerges from interactions between components of all dimensions. Critically, any given component affects all components of lower dimensions on its boundary and all components of higher dimensions containing it. Transport phenomena, including mass diffusion, heat conduction, and charge transport, occur differently through bulk regions, along grain boundaries and junction lines, with each pathway characterised by distinct material properties trovalusci2016MSM , brancherie2017MSM . The evolution of these structures involves discrete, finite changes in arrangement, shape, size, and nature of components through processes such as shear, separation, diffusion, dissolution, and chemical reactions. These microstructural features localise and channel transport, making their accurate representation essential for predictive modelling burczynski2022MSM , baniassadi2023MCM .
From a computational perspective, these multi-scale structures present unique challenges: (i) the need for structure-preserving discretisations that maintain conservation properties exactly, (ii) efficient solution of the resulting algebraic systems with heterogeneous material properties, and (iii) numerical methods that handle arbitrary polytopes without geometric quality constraints. Moreover, the discrete topological changes in material evolution require numerical frameworks that can naturally accommodate discontinuities and interface phenomena without artificial smoothing or enrichment techniques.
@@
Current modelling approaches fail to capture this multi-dimensional complexity adequately. Continuum methods based on smooth topology morro2023Continuum operate with averaged intensive properties and lead to partial differential equations that require numerical approximation. The most popular approaches - finite element methods brenner2008FEM , finite difference methods thomas1995FDM , and finite volume methods leveque2012FVM - effectively capture macroscopic properties but struggle with discontinuities at interfaces and defects. These are addressed through various approximations including cohesive zone models ruiz2001CZM , park2011CZM , extended finite element methods mohammadi2012XFEM , vellwock2024XFEM , and phase-field methods chen2022phase , chen2024phase . However, phase-field approaches use artificial smoothing for sharp transitions, potentially compromising accuracy for localised phenomena. Furthermore, continuum constitutive laws may inadequately capture the complexities of real materials, particularly under extreme deformations, multi-phase interactions, or non-equilibrium processes.
Discrete topology methods offer an alternative foundation through particle-based approaches including molecular dynamics santamaria2023MD , peridynamics madenci2014PD , smoothed particle hydrodynamics filho2019SPH , and discrete element methods jebahi2015DEM . These methods naturally handle atomic and mesoscopic phenomena but lack explicit topological connectivity, making it difficult to enforce constraints and predict emergent behaviours. Their absence of intrinsic volume concepts hampers effective treatment of long-range interactions. A significant limitation is their reliance on heuristic force laws that may lack physical justification.
@@
The constraints of both discrete and continuous topologies can be overcome by adopting cell complexes kozlov2008 , knudson2022 as the topological foundation for modelling. Cell complexes are collections of cells of different topological dimensions arranged following specific rules, studied in algebraic topology. Unlike continuum approaches relying on smooth differentiable manifolds, or discrete particle methods lacking explicit topological structure, cell complexes naturally encode the connectivity and geometry of complex material domains. This makes them particularly suited for modelling materials with intricate microstructures, such as polycrystalline metals, composite materials, and biological tissues.
@@
In previous work berbatov2022diffusion , we demonstrated a formulation of scalar conservation laws (mass, energy, charge, volume) on cell complexes analogous to the strong (differential) form in continuum settings. The formulation used combinatorial differential forms, introduced by Forman forman2002combinatorial , extended with metric-dependent operations for physics applications. We showed construction of boundary value problems on cell complexes and applied the method to diffusion in composites with complex component arrangements. However, the strong form’s analogue approximated Neumann boundary conditions. Weak (integral) formulations avoid this approximation and provide the natural setting for numerical implementation.
1.2 Overview and contributions
In Section 2 we recall the essential elements of exterior calculus on smooth manifolds. The presentation is self-contained, offering abstract definitions of key terms used throughout the paper. While this overview is not a substitute for a textbook on differential geometry (e.g., lee2012introduction ), it introduces the mathematical machinery essential for the developments that follow. It also lays the foundation for the discrete mesh calculus introduced in Section 4, where all operations have natural counterparts in exterior calculus.
In Section 3, we formulate the continuum model for transport phenomena using exterior calculus. Two variational formulations on a $D$-manifold $M$ are derived: a primal weak formulation with the potential as a 0-form (Section 3.1), and a mixed weak formulation with the flow rate as a (D1)(D-1)-form and the dual potential as a $D$-form (Section 3.2). Crucially, all physical quantities are defined on their natural geometric domains as differential forms, and the governing laws are expressed in the language of exterior calculus. To the best of our knowledge, such modelling of transport phenomena has not previously appeared in the literature. While the formulations could be rephrased in vector calculus, their expression in exterior calculus better captures the underlying physical structure. We note that the formulation highlights structural parallels between the continuous and discrete settings, as shown in Section 5, offering conceptual clarity for further theoretical development, and we believe it has significant pedagogical value.
Although the variational derivations conceptually mirror those in vector calculus, we derive them entirely within the language of exterior calculus. The resulting formulations – particularly the mixed one – are related to known results, such as the mixed variational formulation for the Dirichlet problem, a special case of the more general weak Hodge Laplacian problem in Finite Element Exterior Calculus (FEEC) [arnold2018finite, , Chapter 8]. However, our formulation extends the FEEC approach in two key ways: (1) it accommodates mixed boundary conditions (Dirichlet and Neumann simultaneously); (2) while FEEC focuses on the mathematical analysis of Hilbert complexes and their discretisations, we emphasise physical interpretation, discussing the natural domains and physical dimensions of the involved quantities.
In Section 4 we introduce the formalism of combinatorial meshes and their Forman subdivisions. This builds on our previous work in berbatov2022diffusion , which developed meshes of simple polytopes and combinatorial differential forms forman2002combinatorial , interpreted as cochains on subdivisions of quasi-cubical meshes (meshes with cube-like connectivity). Notable differences from berbatov2022diffusion include: (1) we now work with combinatorial meshes – intrinsically defined discrete structures that can be embedded as manifold subdivisions, including with non-flat cells; (2) we adopt the inner product introduced in later work berbatov2023discrete ; (3) we drop the convexity requirement on mesh cells, leading to adjustments in the inner product and Hodge star definitions; and (4) we adopt the standard sign convention for the Hodge star.
In Section 5, we present the primal and mixed variational formulations for transport phenomena on combinatorial meshes. Although this section contains the core contribution of the paper, it is also the most straightforward. With the continuum variational formulations established in Section 3 and the discrete machinery developed in Section 4, the discrete formulations follow by direct translation. The paper is structured to build the theory step by step - from the smooth to the discrete setting - ensuring all necessary mathematical foundations are established before their synthesis.
Finally, in Section 6, we verify our approach with several steady-state problems across different spatial dimensions and with varying domain and mesh geometries. While our long-term goal is to apply the method to complex microstructures (as in [berbatov2022diffusion, , Section 4]), here we test it on manufactured solutions of continuum problems and report discretisation errors. The method can be viewed both as a discretisation of continuum models and as an intrinsically discrete modelling framework. The accompanying code, available at https://github.com/kipiberbatov/cmc, includes additional steady-state and transient examples, though only steady-state results are reported in this article for clarity.
1.3 Style and notations
@@
The sections are broken down into items named discussion, definition, notation, remark, proposition, corollary, formulation, and example, which are numbered to facilitate cross-referencing. All unary operators (excluding taking the additive inverse) have higher precedence than binary operators (excluding taking exponents). For instance, $fv\wedge w$ will mean $(fv)\wedge w$. Further, we record extensively the physical dimensions of the quantities, operators, and equations we discuss. If a quantity $v$ has dimension $X$, we will write $v[X]$ when we want to record the dimension of $v$, and $[[v]]=X$ when we do calculations with dimensions. Physical dimensions will be represented by the SI quantities for mass $\mathsf{M}$, length $\mathsf{L}$, time $\mathsf{T}$, temperature $\theta$, and charge $\mathsf{C}$. (All except the last one are base SI quantities; for historical reasons electric current $\mathsf{I}=\mathsf{C}\mathsf{T}^{-1}$ is used instead of $\mathsf{C}$ as a base quantity, but we prefer charge since it is a conserved property of matter.) If an operator $\varphi$ takes quantities $v_{1}[X_{1}],...,v_{n}[X_{n}]$ and returns a quantity $\varphi(v_{1},...,v_{n})[YX_{1}...X_{n}]$, we will say that $\varphi$ is of dimension $Y$, and write $\varphi[Y]$ or $\varphi]]=Y$. Finally, for equations, if $v[X]$ and $w[X]$, we will emphasise their dimensions in equalities, writing $v=w[X]$.
@@ 2 Overview of exterior calculus
We present only the elements of exterior calculus needed in the rest of the work. Some basic knowledge of general topology is required. Experience with smooth manifolds is also beneficial, although we introduce the terminology we use anyway. A comprehensive book on smooth manifolds is, for instance, lee2012introduction . Exterior calculus offers several advantages over vector calculus: its notation is succinct and dimension-independent, and the underlying mathematical objects are coordinate-free and globally defined. The benefits are:
@@
    A clear distinction between manifold’s intrinsic (non-metric) calculus (Section 2.1) and (additional) calculus depending on a structure such as a Riemannian metric (Section 2.2). This elucidates the distinction between different types of physical relations and the mapping of these relations straightforwardly to their discrete counterparts, for which we build a similar calculus.
@@
    A representation of physical quantities by differential forms ready to be integrated over their respective domains. In contrast, in 33D vector calculus, there are only scalar and vector fields, and it is impossible to differentiate between “scalar” quantities living on 0- and $3$-cells, and “vector” quantities living on $1$-cells and $2$-cells. Terms like pseudo-scalar fields and pseudo-vector fields are often used, but they are related to sign invariance under change of orientation. Because sign invariance is an important notion, we will use the term pseudo-forms.
@@ 2.1 Topological operations
It can be proven [lee2012introduction, , Corollary 3.3] that for any $y\in M$, $T_{y}M$ is a $D$-dimensional real vector space, and $\mathfrak{X}M$ is a $D$-dimensional module over $\mathcal{F}M$. In local coordinates (Cartesian coordinates when $M=\mathbb{R}^{D}$) a vector field $X$ and a tangent vector $\left.X\right|_{y}$ have the representation
@@
The dual space $V^{*}$ of $V$ is defined as $\mathop{\rm Hom}\nolimits(V,R)$, the space of linear functions on $V$. If $V$ is finite dimensional, $\dim V=D$, then also $\dim V^{*}=D$. To a basis $(e_{1},...,e_{D})$ of $V$ corresponds the dual basis $(e^{1},...,e^{D})$ of $V$, so that for $1\leq p,q\leq D$, $e^{p}(e_{q})=\delta^{p}_{q}$, where $\delta^{p}_{q}$ is the Kronecker delta ($1$ if $p=q$ and 0 if $p\neq q$).
[…]
@@
$\left.dx^{i_{1}}\right|_{y}\wedge...\wedge\left.dx^{i_{p}}\right|_{y},\ 1\leq i_{1}<i_{2}<...<i_{p}\leq D.$ (2.15)
@@
where the right hand side is the standard derivative of the real-valued function $f\circ\gamma\colon[0,\delta)\to\mathbb{R}$. We say that $v$ is outward-pointing if v-v is inward pointing.
@@
is outward-pointing. Indeed, the curve $\gamma(t)=(1-t)(\cos\varphi_{0},\sin\varphi_{0})$ witnesses the fact that v-v is inward-pointing.
@@
    $i$ is pairing on covectors: for any $v\in V$ and $\omega\in\Lambda^{1}V^{*}\simeq V^{*}$,
@@
Let $M$ be an oriented smooth manifold of dimension $D$ with boundary and orientation $\mathop{\rm or}\nolimits_{M}$, and $S$ be an oriented (D1)(D-1)-dimensional submanifold of $\partial M$ with orientation $\mathop{\rm or}\nolimits_{S}$. Consider an outward-pointing vector field $X$ on $S$. Since $\dim_{\mathcal{F}M}(\Omega^{D-1}S)=1$, there exists unique smooth function $f\colon S\to\mathbb{R}\setminus\{0\}$ such that
@@
Consider an oriented manifold $M$ with boundary $N$, $D=\dim M$, and let $P$ be an oriented (D2)(D-2)-submanifold of $N$. Locally $P$ divides $N$ into exactly two parts $N^{\prime}$ and $N^{\prime\prime}$. (This is the diamond property – there are exactly two “cells” between $M$ and $P$. A similar property holds for convex polytopes as discussed in [ziegler1995lectures, , Theorem 2.7 (iii)].) Moreover, if $N^{\prime}$ and $N^{\prime\prime}$ have the same orientations as $N$ (the induced boundary orientation), then they will have opposite relative orientations with $P$, i.e., $\varepsilon(N^{\prime},P)=-\varepsilon(N^{\prime\prime},P)$. To allow for possible non-induced orientations on $N^{\prime}$ and $N^{\prime\prime}$, we can write the following equality [berbatov2022diffusion, , Supplementary matrial, Proposition Appendix B.13] which is satisfied in general:
@@
    $\int_{M}d\omega=\int_{\partial M}\mathop{\rm tr}\nolimits_{\partial M}\omega,$ (2.47)
@@
    (base case) integration is function evaluation for 0-forms: for any singleton 0-dimensional manifold $\{x\}$ with orientation $\varepsilon\in\{-1,1\}$, and any $f\colon\{x\}\to\mathbb{R}$,
@@
Let $M$ be an oriented smooth manifold with boundary, $D=\dim M$. An important consequence of the graded Leibniz rule, Equation 2.18, and the Stokes-Cartan theorem, Equation 2.47, is the integration by parts formula: for any $p\in\{0,...,D-1\},\ \omega\in\Omega^{p}_{c}M,\ \eta\in\Omega^{D-p-1}_{c}M$,